OpenStudy (anonymous):

x and y vary inversely, and x = 1/5 when y = 5. What is the value of k? A. 1 25 B. 1 C. 5 D. 25

OpenStudy (whpalmer4):

If \(x\) and \(y\) vary inversely, that means \(y = k/x\). Plug in your known values of \(x\) and \(y\) to find \(k\).

OpenStudy (anonymous):

Damn girl how many more questions you gonna ask?

OpenStudy (anonymous):

y=k*1/x y =k/x

OpenStudy (anonymous):

A lot tf , isn't this what this is for ? hop off if you don't wanna see my questions

OpenStudy (anonymous):

5=k/5 find k 25= k

OpenStudy (whpalmer4):

No, OpenStudy is to help you learn, not just give you answers. For example, if you actually learned how to do this, then you might not just use an incorrect answer.

OpenStudy (anonymous):

Cool :* hop off too

OpenStudy (anonymous):

Or i could just remove you account and ban your ip address from this site, how does that sound?

OpenStudy (anonymous):

5=k/(1/5)

OpenStudy (anonymous):

actually

OpenStudy (anonymous):

Do what you gotta do if that makes you happy then go ahead , I aint trippeeeeeennnn :*

OpenStudy (anonymous):

Just making it so you can LEARN instead of cheat your way through life.

OpenStudy (anonymous):

thanks ^ you really are a big help boo :*